OpenStudy (anonymous):

Write the equation of the line with a slope of -1/3 and passing through the point (6, -4).

Parth (parthkohli):

Nawww, point slope.

Parth (parthkohli):

\[\rm y - y_1 = m(x - x_1)\]

Parth (parthkohli):

Substitute!\[\rm y - (-4) = {-1 \over 3}\left(x - 6 \right) \]...\[\rm y + 4 = -{x \over 3} +2\]

Parth (parthkohli):

Done :D

OpenStudy (anonymous):

thank you so much!

Parth (parthkohli):

No, it's not done @stolas

Parth (parthkohli):

I was just playing a little.

Parth (parthkohli):

Hello? @stolas